A container holds 0.70 g of oxygen at a pressure of 8.3 atm .

a. How much heat is required to increase the temperature by 119 ∘C at constant pressure?

b.How much will the temperature increase if this amount of heat energy is transferred to the gas at constant volume?
Answer #1

number of moles

n = given mass / molar mass = 0.7 / 32

n = 0.022 moles

enervy needed for diatomic gas

q = (7/2) n R (T2 - T1)

q = 3.5* 0.022* 8.314* 119

q = 75.748

======

b)

for constant volume diatomic gas

q = (5/2) nR (T2 - T1)

75.748 = 2.5* 0.022* 8.314* dT

dT = 165.65 C

======